OpenStudy (anonymous):

4(5)-4/2 multiplied by 6... help please!

Parth (parthkohli):

=20 - 2 = 18

OpenStudy (callisto):

Do you mean \[\frac{4(5)-4}{2}\times 6\]?

Parth (parthkohli):

18 * 6 = 108

Parth (parthkohli):

108 is the answer.

OpenStudy (anonymous):

20-2=18 28*6= 108 :)